# On the Role of Attention Masks and LayerNorm in Transformers

### v1

@@ 6 Conclusion
The attention mechanism has led to significant empirical progress in building foundation models. From a dynamical system perspective, the time-varying and nonlinear nature of self-attention dynamics, when coupled with attention masks and LayerNorm, enables expressive and complex behavior. In particular, we show that the choice of attention masks directly affects the token dynamics. It would hence be valuable for future research to investigate how to effectively design attention masks. Our result further suggests that robust token geometries under multi-layer self-attention can exhibit both full-rankness and anisotropic characteristics simultaneously, which is the case in real transformers as well. It would be interesting to study how such co-existence of these two characteristics would help with learning tasks — whether the full-rank yet close-to-low-dimensional geometry enables efficient learning and generalization while still letting tokens capture meaningful fine-grained details.
@@ Acknowledgments
X.W., A.A., and A.J. would like to thank Bernard Chazelle for helpful discussions. This research has been supported in part by ONR Award N00014-23-1-2299 and a Vannevar Bush fellowship from the Office of the Under Secretary of Defense for Research and Engineering (USD(R&E)).
@@ B.1 Auxiliary results
Let $\{A^{(n)}\}_{n=0}^{\infty}$ in $\mathcal{A}_{\mathcal{G},\epsilon}$ and $r$ be the diameter of $\mathcal{G}$. Then for $t\geq 0$ and $0\leq p\leq 1$,
@@
Iterating this process for tokens in $\mathcal{V}_{3},...,\mathcal{V}_{r}$, where $r$ is the diameter of $\mathcal{G}$, we get that for any token $i\in\mathcal{G}$,
@@
where $r$ is the diameter of the graph $\mathcal{G}$.
@@ C.1 Auxiliary results
First, since for $x\in\operatorname{Conv}(X^{(t)})$, ${(W^{(t)})}^{\top}x$ always lies in $\mathbb{S}^{d-1}$, $D^{(t)}_{i,i}\geq 1$.
@@
Given by assumption, all the points of $X^{(0)}$ lie in an open hemisphere $\mathcal{H}$. Since all the points of $X^{(0)}$ lie in an open hemisphere $\mathcal{H}$, there exists $\gamma>0$ such that for all $x\in\operatorname{Conv}(X^{(0)})$ and $W$ orthogonal,
@@ C.1.1 Exponential convergence rate
$\phi^{(t+r)}$ $=\underset{i,j\in[N]}{\min}\langle X^{(t+r)}_{i,:},X^{(t+r)}_{i,:}\rangle=\langle X^{(t+r)}_{i^{(t+r)},:},X^{(t+r)}_{j^{(t+r)},:}\rangle$
@@ Appendix D Proof of Corollary 1
First, since for $x\in\operatorname{Conv}(X^{(t)})$, ${(W^{(t)})}^{\top}x$ always lies in $\mathbb{S}^{d-1}$, $D^{(t)}_{i,i}\geq 1$, for all $i\in[N],t\geq 0$.
@@
$\phi^{(t+r)}$ $=\underset{i,j\in[N]}{\min}\langle X^{(t+r)}_{i,:},X^{(t+r)}_{i,:}\rangle=\langle X^{(t+r)}_{i^{(t+r)},:},X^{(t+r)}_{j^{(t+r)},:}\rangle$
@@
$\geq\sum_{(k_{1},...,k_{r})\in[N]^{d}}\sum_{(l_{1},...,l_{r})\in[N]^{d}}A^{(t+r-1)}_{i^{(t+r)},k_{1}}...A^{(t)}_{k_{r-1},k_{r}}A^{(t+r-1)}_{j^{(t+r)},l_{1}}...A^{(t)}_{l_{r-1},l_{r}}\langle X^{(t)}_{k_{r}},X^{(t)}_{l_{r}}\rangle$
$\geq N\epsilon^{2r}+(1-N\epsilon^{2r})\phi^{(t)}\,.$
@@
$1-\phi^{(kr)}\leq(1-N\epsilon^{2r})^{k}(1-\phi^{(0)})\,,$
Let $C\vcentcolon=1/(1-N\epsilon^{2r})$. Hence for all $t\geq 0$,
$1-\phi^{(t)}\leq C(1-N\epsilon^{2r})^{\frac{t}{r}}\,.$ (27)
@@
$\leq C^{\prime}(1-N\epsilon^{2r})^{\frac{t}{2r}}\,,$
@@ G.3 The evolution of singular values of token representations in transformers
In this section, we investigate the full set of singular values of token representations $X^{(t)}$ in transformers. For a clear presentation, we randomly select $3$ out of the $32$ sequence samples described in Section 5 and calculate all the singular values of $X^{(t)}$. The results in randomly initialized and pretrained transformers are presented in Fig. 8.
We see that in both pretrained and initialized transformers, while the rank is full for $X^{(t)}$, there is usually one principal component that is much more dominant than the rest. As we have discussed in the main text, this is the anisotropic characteristic of token representations— there is a strong non-uniformity in different directions. Nonetheless, it is important to note that the “small" singular values are in fact not negligible here in the absolute sense—the minimal singular values are never close to zero, as Fig. 4 (middle) shows.
@@ G.4 Initial token geometry
We verify the initial token geometry in transformers. For a clear presentation, we randomly select $3$ out of the $32$ sequence samples described in Section 5 and calculate all the pairwise cosine similarities between tokens. The results in randomly initialized and pretrained transformers are presented in Fig. 9. We see that in particular in BERT and ALBERT, the initial pairwise cosine similarities are all non-negative.

### v2

@@ 6 Conclusion
The attention mechanism has led to significant empirical progress in building foundation models. From a dynamical system perspective, the time-varying and nonlinear nature of self-attention dynamics, when coupled with attention masks, LayerNorm and other components in transformers, enables expressive and complex behavior. In particular, we show that the choice of attention masks directly affects the token dynamics. It would hence be valuable for future research to investigate how to effectively design attention masks. Our result further suggests that robust token geometries under multi-layer self-attention can exhibit both full-rankness and anisotropic characteristics simultaneously, which is the case in real transformers as well. It would be interesting to study how such co-existence of these two characteristics would help with learning tasks — whether the full-rank yet close-to-low-dimensional geometry enables efficient learning and generalization while still letting tokens capture meaningful fine-grained details.
@@ Acknowledgments
X.W., A.A., and A.J. would like to thank Bernard Chazelle for helpful discussions. X.W., A.A., and A.J. were supported by ONR Award N00014-23-1-2299 and a Vannevar Bush fellowship from the Office of the Under Secretary of Defense for Research and Engineering (USD(R&E)). Y.W. and S.J. were supported by ONR Award N00014-20-1-2023 (MURI ML-SCOPE), NSF AI Institute TILOS (NSF CCF-2112665), NSF Award 2134108, and the Alexander von Humboldt Foundation.
@@ B.1 Auxiliary results
Let $\{A^{(n)}\}_{n=0}^{\infty}$ in $\mathcal{A}_{\mathcal{G},\epsilon}$ and $r$ be the radius of $\mathcal{G}$. Then for $t\geq 0$ and $0\leq p\leq 1$,
@@
Iterating this process for tokens in $\mathcal{V}_{3},...,\mathcal{V}_{r}$, where $r$ is the radius of $\mathcal{G}$, we get that for any token $i\in\mathcal{G}$,
@@
where $r$ is the radius of the graph $\mathcal{G}$.
@@ C.1 Auxiliary results
First, since for $x\in\operatorname{Conv}(X^{(t)})$, ${W^{(t)}}^{\top}x$ always lies in $\mathbb{S}^{d-1}$, $D^{(t)}_{i,i}\geq 1$.
@@
We then make the following observation:
Lemma 6.
Under assumption (*), there exists $v\neq\mathbf{0}$ such that
$\langle X^{(0)}_{i,:},v\rangle>0\quad i\in[N]\,,$
meaning that all the points of $X^{(0)}$ lie in an open hemisphere $\mathcal{H}$.
To see this, observe that since $X^{(0)}$ is full rank and $N\leq d$, $\mathbf{0}\notin\operatorname{Conv}(X^{(0)})$. Then the hyperplane separation theorem states that there exists a hyperplane strictly separating $\mathbf{0}$ and $\operatorname{Conv}(X^{(0)})$, letting us conclude Lemma 6.
Since all the points of $X^{(0)}$ lie in an open hemisphere $\mathcal{H}$, there exists $\gamma>0$ such that for all $x\in\operatorname{Conv}(X^{(0)})$ and $W$ orthogonal,
@@ C.1.1 Exponential convergence rate
$\phi^{(t+r)}$ $=\underset{i,j\in[N]}{\min}\langle X^{(t+r)}_{i,:},X^{(t+r)}_{j,:}\rangle=\langle X^{(t+r)}_{i^{(t+r)},:},X^{(t+r)}_{j^{(t+r)},:}\rangle$
@@ Appendix D Proof of Corollary 1
First, since for $x\in\operatorname{Conv}(X^{(t)})$, ${W^{(t)}}^{\top}x$ always lies in $\mathbb{S}^{d-1}$, $D^{(t)}_{i,i}\geq 1$, for all $i\in[N],t\geq 0$.
@@
$\phi^{(t+r)}$ $=\underset{i,j\in[N]}{\min}\langle X^{(t+r)}_{i,:},X^{(t+r)}_{j,:}\rangle=\langle X^{(t+r)}_{i^{(t+r)},:},X^{(t+r)}_{j^{(t+r)},:}\rangle$
@@
$\geq\sum_{(k_{1},...,k_{r})\in[N]^{r}}\sum_{(l_{1},...,l_{r})\in[N]^{r}}A^{(t+r-1)}_{i^{(t+r)},k_{1}}...A^{(t)}_{k_{r-1},k_{r}}A^{(t+r-1)}_{j^{(t+r)},l_{1}}...A^{(t)}_{l_{r-1},l_{r}}\langle X^{(t)}_{k_{r}},X^{(t)}_{l_{r}}\rangle$
$\geq\epsilon^{2r}+(1-\epsilon^{2r})\phi^{(t)}\,.$
@@
$1-\phi^{(kr)}\leq(1-\epsilon^{2r})^{k}(1-\phi^{(0)})\,,$
Let $C\vcentcolon=1/(1-\epsilon^{2r})$. Hence for all $t\geq 0$,
$1-\phi^{(t)}\leq C(1-\epsilon^{2r})^{\frac{t}{r}}\,.$ (28)
@@
$\leq C^{\prime}(1-\epsilon^{2r})^{\frac{t}{2r}}\,,$
@@ G.3 The evolution of singular values of token representations in transformers
In this section, we investigate the full set of singular values of token representations $X^{(t)}$ in transformers. For a clear presentation, we randomly select $3$ out of the $3000$ sequence samples described in Section 5 and calculate all the singular values of $X^{(t)}$. The results in randomly initialized transformers are presented in Fig. 8.
We see that while the rank is full for $X^{(t)}$, there is usually one principal component that is much more dominant than the rest. As we have discussed in the main text, this is the anisotropic characteristic of token representations— there is a strong non-uniformity in different directions. Nonetheless, it is important to note that the “small" singular values are in fact not negligible here in the absolute sense—the minimal singular values are never close to zero, as Fig. 4 (middle) shows.
@@ G.4 Initial token geometry
We verify the initial token geometry in transformers. For a clear presentation, we randomly select $3$ out of the $3000$ sequence samples described in Section 5 and calculate all the pairwise cosine similarities between tokens. The results in randomly initialized transformers are presented in Fig. 9. We see that in particular in BERT and ALBERT, the initial pairwise cosine similarities are all non-negative.